# Transit Campus, Government ITI, Berhampur 760010, Odisha, India

### v1

@@ 4.2.1 $\chi$-Related Parameters
From the left panel of the figure, it can be seen that at fixed $g_{B-L}$, extremely small change in the value of $y_{1,2,3}$ can drastically change the value of abundance. The explicit dependence is very hard to comprehend from the left panel of the figure. Therefore, in order to show the strong dependence of abundance on the couplings, we take a very small parameter range on the right panel of Fig. 5. It can be seen from the right panel that at fixed $g_{B-L}$ values, $f_{\mathrm{PBH}}$ increases with decreasing $y_{1,2,3}$. Due to this robust dependence, abundance of PBH can be used as a tool to put bounds on such couplings. Next, we focus on the mass of the PBH created during the FOPT. We show the relation of the mass of the PBH and the couplings in Fig. 6 keeping $v_{\chi}=10^{8}\mathrm{~{}GeV}$. It is evident from the figure that the mass of PBH increases with decreasing $g_{B-L}$, but the dependence of $M_{\mathrm{PBH}}$ on $y_{1,2,3}$ is very weak.
Figure 6: We show the dependence of the mass of the PBH generated from the FOPT on the couplings $g_{B-L}$ and $y_{1,2,3}$. In this case $v_{\chi}=10^{8}\mathrm{~{}GeV}$.
It is to be noted here, though we have not shown the explicit dependence of the PBH abundance on the VEV, it can be inferred from Figs. 42, and 3.
Now we focus on the SGWB due to the FOPT driven by $\chi$. As mentioned before, the spectrum and the peak frequency of the SGWB depend on the FOPT parameters, i.e., $\alpha$, $\beta/H$, $T_{p}$ and those in turn depend on $g_{B-L}$, $y_{1,2,3}$, and $v_{\chi}$. From the above figures, we have seen that $g_{B-L}$ has the strongest effect on $\alpha$ and $\beta/H$ whereas, $v_{\chi}$ has the strongest dependence on the percolation temperature. Therefore, in Fig. 7 we show the dependence on the SGWB spectrum arising from the FOPT on $g_{B-L}$ (left panel) and $v_{\chi}$ (right panel) respectively.
Figure 7: (Left) We show the SGWB spectra for varying $g_{B-L}$ while keeping $y_{1,2,3}=0.1$ and $v_{\chi}=10^{8}\mathrm{~{}GeV}$. (Right) We show the SGWB spectra for varying $v_{\chi}$ while keeping $y_{1,2,3}=0.2$ and $g_{B-L}=0.1$. In both cases we show the sensitivity of the proposed GW detectors such as Taiji Ruan:2018tsw , LISA LISA:2017pwj , LVK Somiya:2011np ; LIGOScientific:2014pky , CE Reitze:2019iox , ET Punturo:2010zz , DECIGO Kawamura:2011zz , and BBO Phinney:2004bbo .
It is evident that at higher values of $g_{B-L}$ the peak frequency increases and the amplitude decreases. Whereas, for increasing values of $v_{\chi}$ the peak frequency of the GW spectra increases while the amplitude remains unchanged.
@@ 4.2.2 $\Phi_{1,2}$-Related Parameters
Here we focus on the possibility of a second FOPT, i.e., the EWSB driven by the doublets. It is to be noted here in this regard, Ref. Benincasa:2022elt has shown FOPTs driven by both $\Phi_{1}$ and $\Phi_{2}$. However, the FOPTs which may occur in the physically allowed parameter space are much weaker than the one driven by $\chi$ and hence for simplicity we constrain ourselves in to the FOPTs driven by SM higgs with contributions from the BSM doublet.
@@
$\lambda_{1}$ $=\dfrac{m_{h}^{2}}{2v^{2}}$ $\lambda_{7}$ $=-\dfrac{m_{h}^{2}}{2v_{\chi}^{2}}$ $\lambda_{8}$ $=\dfrac{2m_{H}^{2}-\lambda_{356}v^{2}}{2v_{\chi}^{2}}$
$\lambda_{3}$ $=\lambda_{356}+2\dfrac{m_{H^{\pm}}^{2}-m_{H}^{2}}{v^{2}}$ $\lambda_{5}$ $=\dfrac{m_{H}^{2}+m_{A}^{2}-2m_{H^{\pm}}^{2}}{v^{2}}$ $\lambda_{6}$ $=\dfrac{m_{H}^{2}-m_{A}^{2}}{v^{2}}$ (23)
where $m_{h,H,A,H^{\pm}}$ are the masses of the SM higgs, the extra neutral scalars $H$ and $A$ and the charged scalar $H^{\pm}$ and $\lambda_{356}=\lambda_{3}+\lambda_{5}+\lambda_{6}$. It is worth mentioning here that $H$($A$) will play the role of particle DM if the the mass of $H$ is lower (higher) than $A$. From the above relations we can see that since the mass of the SM higgs is well measured the coupling $\lambda_{7}$ depends on $v_{\chi}$. For example, SM higgs mass can be obtained if $v_{\chi}=10^{7}\mathrm{~{}GeV}$ and $\lambda_{7}=-7.8125\times 10^{-11}$. Therefore, in order to facilitate the EWSB in our model, the coupling $\lambda_{7}$ has to be specifically fine-tuned. Although $\lambda_{8}$ has less constraints on it due to $m_{H}$ and $\lambda_{356}$ being variables, still its order will be determined by $v_{\chi}$ due to its value being much larger than $v$ which has been experimentally determined to be 246 GeV. The values of $\lambda_{3}$, $\lambda_{5}$, and $\lambda_{6}$ are determined by $m_{H,A,H^{\pm}}$ along with the equality that $\lambda_{356}=\lambda_{3}+\lambda_{5}+\lambda_{6}$.
It has been shown that the mass of the DM candidate has to be larger than $\mathcal{O}(60)\mathrm{~{}GeV}$ for FOPTs to happen in the single step configuration Benincasa:2022elt . Moreover, to satisfy the relic density constraint one has to resort to only the resonance region i.e., $\sim 62.5$ GeV. Since in our model we can have PBH as the primary DM candidate we relax the constricted mass range for the scalar DM and take $m_{H}\in[200,1000]\mathrm{~{}GeV}$. Furthermore, it is also shown that $\lambda_{2}$ also does not have a strong dependence on FOPT. However, $\lambda_{356}$ has a strong dependence on the possibility of a FOPT. Therefore, remaining within the parameter space verified in Ref. Benincasa:2022elt , we show the dependence of $\lambda_{356}$ and $m_{H}$ on the FOPT parameters for completeness.
@@
It can be seen from Fig. 8, for the entire parameter space, the $\alpha$ and $\beta/H$ are very small and very large respectively. Furthermore, it can also be seen that both $\alpha$ and $\beta/H$ has much stronger dependence on $m_{H}$ as compared to $\lambda_{356}$. It is worth mentioning here that for this FOPT, we find $T_{n}\sim T_{p}$. For the entire parameter space that we consider, $T_{n}\in[156,199]\mathrm{~{}GeV}$.
@@
In Fig. 9 we show the dependence of the SGWB spectrum originating from this FOPT on $m_{H}$ and $\lambda_{356}$. It can be seen from the figure that as $\lambda_{356}$ increases and $m_{H}$ decreases the amplitude of the SGWB spectra increases. In our parameter space for $\lambda_{356}=10$ and $m_{H}=200\mathrm{~{}GeV}$ we obtain the strongest SGWB signal which can be detected in DECIGO and BBO. However, it should be noted that the $\chi$ driven FOPT creates much larger SGWB, which will overshadow the SGWB due to the $\Phi_{1,2}$ driven FOPT.
@@ 4.3 Benchmark Parameters and Consequences
In Tab. 3 we show the consequences of the benchmark cases considered in Tab. 2. It can be seen that in both the cases, PBHs effectively contributed to the entire DM fraction of the universe. Furthermore, each of these benchmark cases allows for two FOPTs at different temperatures, hence contributes to two different SGWB spectra which we show in Fig. 10. It can be seen that for each case, the FOPT driven by $\chi$ creates SGWB spectra which is in the sensitivity range of LISA, Taiji, DECIGO, BBO, CE, and ET. However, in both the cases the SGWB created during the FOPT driven by the doublets are out of the sensitivity range of the detectors. It is worth mentioning that in both these cases the relic density of the particle DM is $\mathcal{O}(10^{-4})$ and PBH plays the role of primary DM in the universe.
BP $f_{\mathrm{PBH}}$ $M_{\mathrm{PBH}}(\mathrm{g})$ $\mathrm{PT}_{\chi}(\alpha,~{}\beta/H,~{}T_{p}(\mathrm{~{}GeV}))$ $\mathrm{PT}_{\Phi_{1,2}}(\alpha,~{}\beta/H,~{}T_{n}(\mathrm{~{}GeV}))$
1 0.9997 $1.25\times 10^{21}$ $(13200,~{}26.8,~{}2.71\times 10^{5})$ $(0.013,~{}3905,~{}161)$
2 0.999 $4.83\times 10^{22}$ $(2700,~{}26.7,~{}4.73\times 10^{4})$ $(0.0088,~{}4850,~{}176.2)$
@@ 5 Conclusion
In order to quantify the properties of the FOPTs we scan different parameter spaces. We find that for the high temperature FOPT, the strength ($\alpha$) decreases sharply with the $U(1)_{B-L}$ gauge coupling $g_{B-L}$. However, $\alpha$ has a very weak dependence on the Yukawa couplings $y_{1,2,3}$ and the $v_{\chi}$. The inverse duration of the FOPT, $\beta/H$ increases sharply with $g_{B-L}$, and similar to the previous case, it has very weak dependence on $y_{1,2,3}$ and $v_{\chi}$. We further find that the critical, nucleation, and percolation temperatures increase with $g_{B-L}$ and $v_{\chi}$ and they have very weak dependence on $y_{1,2,3}$. For the case of PBH, we found that the abundance is extremely sensitive to the gauge and Yukawa couplings, i.e., for very small increase of the couplings, the abundance decreases drastically. Furthermore, in the coupling range where formation of PBH is allowed, the PBH mass decreases with increasing coupling. The parameter space we consider, generates SGWB spectra is at the detectable range of the future experiments LISA, Taiji, DECIGO, BBO, ET, CE, etc. For the second phase transition we just vary the mass of $H$ and $\lambda_{356}$ to see their effect on the FOPT parameters. We find that $\alpha$ ($\beta/H$) decreases (increases) with increasing mass of $H$, and both of them has very weak dependence on $\lambda_{356}$. We also see that for a very small part of the parameter range there can be detectable SGWB spectra. We see that the amplitude of the SGWB increases with decreasing $m_{H}$ and increasing $\lambda_{356}$.
@@
Next, we consider two sets of benchmark values and see their consequences in the PBH and SGWB aspect. We find that though in both the cases there is negligible relic density of the particle DM, the PBH plays the role of entire DM. We also find that in both the cases the SGWB only from the $\chi$ driven FOPT is detectable.
@@
It is worth mentioning here that there are a few conditions and approximations that we have used throughout the article. The first one is that the strong $\chi$ driven phase transition only happens when the self-interaction of $\chi$ is negligible. We also have assumed equal Yukawa couplings between the three right handed neutrinos and the singlet scalar. Though for our purpose the unequal couplings would not have changed the results, it can still have other phenomenological implications which we do not focus on. Finally, it is to be noted that different prescriptions available in different literature regarding the creation of PBH during a FOPT and its abundance suggests different values of upper bound of $\beta/H$ for which PBHs can be created. However, we follow one of them since our motivation is to show that various models which have tight constraints on them from the DM aspect, can be revived by using a $U(1)$ extension which can give us FOPTs strong and slow enough to create PBHs with appreciable abundance. In conclusion, we claim that these kind of extensions can be applied to other models as well which have rich phenomenological implications but is heavily constrained from the DM relic density aspect.

### v2

@@ 4.2.1 $\chi$-Related Parameters
From the left panel of the figure, it can be seen that at fixed $g_{B-L}$, extremely small change in the value of $y_{1,2,3}$ can drastically change the value of abundance. The explicit dependence is very hard to comprehend from the left panel of the figure. Therefore, in order to show the strong dependence of abundance on the couplings, we take a very small parameter range on the right panel of Fig. 4. It can be seen from the right panel that at fixed $g_{B-L}$ values, $f_{\mathrm{PBH}}$ increases with decreasing $y_{1,2,3}$. Due to this robust dependence, abundance of PBH can be used as a tool to put bounds on such couplings. Next, we focus on the peak mass of the PBH created during the FOPT. We show the relation of the mass of the PBH and the gauge coupling and the VEV in Fig. 5 keeping the Yuakwa couplings $y_{1,2,3}=0.2$. It is evident from the figure that the mass of PBH increases with decreasing $g_{B-L}$ and $v_{\chi}$.
Figure 5: We show the dependence of the mass of the PBH generated from the FOPT on the coupling $g_{B-L}$ and the VEV $v_{\chi}$. In this case $y_{1,2,3}=0.2$.
Now we focus on the SGWB due to the FOPT driven by $\chi$ for which, as mentioned before, we consider the contributions from the collision of the bubble walls and the curvature perturbation. As mentioned before, the spectrum and the peak frequency of the SGWB depend on the FOPT parameters, i.e., $\alpha$, $\beta/H$, $T_{\mathrm{reh}}$ and those in turn depend on $g_{B-L}$, $y_{1,2,3}$, and $v_{\chi}$. Therefore, in Fig. 6 we show the dependence on the SGWB spectrum arising from the FOPT on $g_{B-L}$ (top-left panel), $y_{1,2,3}$ (top-right panel), and $v_{\chi}$ (bottom panel) respectively.
Figure 6: (Top-left) We show the SGWB spectra for varying $g_{B-L}$ while keeping $y_{1,2,3}=0.18$ and $v_{\chi}=10^{8}\mathrm{~{}GeV}$. (Top-right) We show the SGWB spectra for varying $y_{1,2,3}$ while keeping $g_{B-L}=0.2$ and $v_{\chi}=10^{8}\mathrm{~{}GeV}$. (Bottom) We show the SGWB spectra for varying $v_{\chi}$ while keeping $y_{1,2,3}=0.27$ and $g_{B-L}=0.25$. In both cases we show the sensitivity of the proposed GW detectors such as Taiji Ruan:2018tsw , LISA LISA:2017pwj , LVK Somiya:2011np ; LIGOScientific:2014pky , CE Reitze:2019iox , ET Punturo:2010zz , DECIGO Kawamura:2011zz , and BBO Phinney:2004bbo .
It is evident that at higher values of $g_{B-L}$ the peak frequency increases and the amplitude decreases. Whereas, for increasing $y_{1,2,3}$ the amplitude of the GW spectra increases but the peak frequency remains constant. Finally, for increasing values of $v_{\chi}$ the peak frequency and the amplitude both increases.
@@ 4.2.2 $\Phi_{1,2}$-Related Parameters
In the context of inert doublet model Ref. Benincasa:2022elt has shown FOPTs driven by both $\Phi_{1}$ and $\Phi_{2}$. Here we focus on the possibility of a second FOPT, i.e., the EWSB driven by the doublets. However, the FOPTs which may occur in the physically allowed parameter space are much weaker than the one driven by $\chi$ and hence for simplicity we constrain ourselves in to the FOPTs driven by SM higgs with contributions from the BSM doublet.
@@
$\lambda_{1}$ $=\dfrac{m_{h}^{2}}{2v^{2}},~{}~{}\lambda_{7}=-\dfrac{m_{h}^{2}}{2v_{\chi}^{2}},~{}~{}\lambda_{8}=\dfrac{2m_{H}^{2}-\lambda_{356}v^{2}}{2v_{\chi}^{2}},$
$\lambda_{3}$ $=\lambda_{356}+\frac{2}{v^{2}}\left(m_{H^{\pm}}^{2}-m_{H}^{2}\right),~{}~{}\lambda_{5}=\frac{1}{v^{2}}\left(m_{H}^{2}+m_{A}^{2}-2m_{H^{\pm}}^{2}\right),~{}~{}\lambda_{6}=\frac{1}{v^{2}}\left(m_{H}^{2}-m_{A}^{2}\right),$ (34)
where $m_{h,H,A,H^{\pm}}$ are the masses of the SM higgs, the extra neutral scalars $H$ and $A$ and the charged scalar $H^{\pm}$ and $\lambda_{356}=\lambda_{3}+\lambda_{5}+\lambda_{6}$. It is worth mentioning here that $H$($A$) will play the role of particle DM if the the mass of $H$ is lower (higher) than $A$. For our parameter choices we find $m_{H}<m_{A}$, hence $H$ is the particle DM candidate. From the above relations we can see that since the mass of the SM higgs is well-measured the coupling $\lambda_{7}$ depends on $v_{\chi}$. For example, SM higgs mass can be obtained if $v_{\chi}=10^{7}\mathrm{~{}GeV}$ and $\lambda_{7}=-7.8125\times 10^{-11}$. Therefore, in order to facilitate the EWSB in our model, the coupling $\lambda_{7}$ has to be specifically fine-tuned. Although $\lambda_{8}$ has less constraints on it due to $m_{H}$ and $\lambda_{356}$ being variables, still its order will be determined by $v_{\chi}$ due to its value being much larger than $v$ which has been experimentally determined to be 246 GeV. The values of $\lambda_{3}$, $\lambda_{5}$, and $\lambda_{6}$ are determined by $m_{H,A,H^{\pm}}$ along with the equality that $\lambda_{356}=\lambda_{3}+\lambda_{5}+\lambda_{6}$.
@@
It has been shown that the mass of the DM candidate has to be larger than $\mathcal{O}(60)\mathrm{~{}GeV}$ for FOPTs to happen in the single step configuration Benincasa:2022elt . Moreover, to satisfy the relic density constraint one has to resort to only the resonance region i.e., $\sim 62.5$ GeV. Since in our model we can have PBH as the primary DM candidate we relax the constricted mass range for the scalar DM and take $m_{H}\in[200,1000]\mathrm{~{}GeV}$. Furthermore, it is also shown that $\lambda_{2}$ also does not have a strong dependence on FOPT. However, $\lambda_{356}$ has a strong dependence on the possibility of a FOPT. Therefore, remaining within the parameter space verified in Ref. Benincasa:2022elt , we show the dependence of $\lambda_{356}$ and $m_{H}$ on the FOPT parameters for completeness. It can be seen from Fig. 7, for the entire parameter space, the $\alpha$ (left panel) and $\beta/H$ (right panel) are very small and very large respectively. Furthermore, it can also be seen that both $\alpha$ and $\beta/H$ has much stronger dependence on $m_{H}$ as compared to $\lambda_{356}$. It is worth mentioning here that for this FOPT, we find $T_{n}\sim T_{p}$. For the entire parameter space that we consider, $T_{n}\in[156,199]\mathrm{~{}GeV}$.
@@
In Fig. 8 we show the dependence of the SGWB spectrum originating from this FOPT on $m_{H}$ (left panel) and $\lambda_{356}$ (right panel). It can be seen from the figure that as $\lambda_{356}$ increases and $m_{H}$ decreases the amplitude of the SGWB spectra increases. In our parameter space for $\lambda_{356}=10$ and $m_{H}=200\mathrm{~{}GeV}$ we obtain the strongest SGWB signal which can be detected in DECIGO and BBO. However, it should be noted that the $\chi$ driven FOPT creates much larger SGWB, which will overshadow the SGWB due to the $\Phi_{1,2}$ driven FOPT.
@@ 4.3 Benchmark Parameters and Consequences
In Tab. 3 we show the consequences of the benchmark cases considered in Tab. 2. It can be seen that in both the cases, PBHs effectively contributed to the entire DM fraction of the universe. Furthermore, each of these benchmark cases allows for two FOPTs at different temperatures, hence contributes to two different SGWB spectra which we show in Fig. 9. It is to be noted that for the SGWB due to the $\chi$ driven FOPT, the left peak is a result of the curvature perturbation whereas the right peak is a result of the collision of bubble walls. It can be seen that for each case, the FOPT driven by $\chi$ creates SGWB spectra which is in the sensitivity range of LISA, Taiji, DECIGO, BBO, CE, and ET. However, in both the cases the SGWB created during the FOPT driven by the doublets are only within the sensitivity ranges of the detectors BBO and DECIGO. It is worth mentioning that in both these cases the relic density of the particle DM is $\mathcal{O}(10^{-4})$ and PBH plays the role of primary DM in the universe.
BP $f_{\mathrm{PBH}}$ $M_{\mathrm{PBH}}(\mathrm{g})$ $\mathrm{PT}_{\chi}(\alpha,~{}\beta/H,~{}T_{\mathrm{reh}}~{}(\mathrm{GeV}))$ $\mathrm{PT}_{\Phi_{1,2}}(\alpha,~{}\beta/H,~{}T_{n}~{}(\mathrm{GeV}))$
1 0.9997 $7.98\times 10^{18}$ $(289,~{}7.65,~{}3.66\times 10^{6})$ $(0.013,~{}3905,~{}161)$
2 0.9999 $6.885\times 10^{20}$ $(80908,~{}7.56,~{}3.94\times 10^{5})$ $(0.0088,~{}4850,~{}176.2)$
@@ 5 Conclusion
In order to quantify the properties of the FOPTs we scan different parameter spaces. We find that for the high temperature FOPT, the strength ($\alpha$) decreases sharply with the $U(1)_{B-L}$ gauge coupling $g_{B-L}$. However, $\alpha$ has weaker dependence on the Yukawa couplings $y_{1,2,3}$ and the $v_{\chi}$. The inverse duration of the FOPT, $\beta/H$ increases sharply with $g_{B-L}$, and similar to the previous case, it has weaker dependence on $y_{1,2,3}$ and $v_{\chi}$. We further find that the critical, nucleation, and percolation temperatures increase with $g_{B-L}$ and $v_{\chi}$ and they have a weaker dependence on $y_{1,2,3}$. For the case of PBH, we found that the abundance is extremely sensitive to the gauge and Yukawa couplings, i.e., for very small increase of the couplings, the abundance decreases drastically. Furthermore, in the coupling range where formation of PBH is allowed, the PBH mass decreases with increasing coupling. The parameter space we consider, generates SGWB spectra is at the detectable range of the future experiments LISA, Taiji, DECIGO, BBO, ET, CE, etc. For the second phase transition we just vary the mass of $H$ and $\lambda_{356}$ to see their effect on the FOPT parameters. We find that $\alpha$ ($\beta/H$) decreases (increases) with increasing mass of $H$, and both of them has very weak dependence on $\lambda_{356}$. We also see that only for a fraction of the parameter range there can be detectable SGWB spectra. We see that the amplitude of the SGWB increases with decreasing $m_{H}$ and increasing $\lambda_{356}$.
@@
Next, we consider two sets of benchmark values and see their consequences in the PBH and SGWB aspect. We find that though in both the cases there is negligible relic density of the particle DM, the PBH plays the role of entire DM. We also find that although in both cases SGWB from both the FOPTs can be detected by future GW detectors, in both the cases the SGWB from the $\chi$ driven FOPT completely overshadows the other one.
@@
It is worth mentioning here that there are a few conditions and approximations that we have used throughout the article. The first one is that the strong $\chi$ driven phase transition only happens when the self-interaction of $\chi$ is negligible. We also have assumed equal Yukawa couplings between the three right handed neutrinos and the singlet scalar. Though for our purpose the unequal couplings would not have changed the results, it can still have other phenomenological implications which we do not focus on. Finally, it is to be noted that different prescriptions available in different literature regarding the creation of PBH during a FOPT and its abundance suggests different values of upper bound of $\beta/H$ for which PBHs can be created. However, we follow one of them since our motivation is to show that various models which have tight constraints on them from the DM aspect, can be revived by using a $U(1)$ extension which can give us FOPTs strong and slow enough to create PBHs with appreciable abundance. In conclusion, we claim that these kind of extensions can be applied to other models as well which have rich phenomenological implications but are heavily constrained from the DM relic density aspect.